OpenStudy (anonymous):

Write an equation for the line that passes through (2,4) and (-1,4).

OpenStudy (anonymous):

find its slope m by (y - y)/(x - x) then substitute y = m x + b, find b then done

OpenStudy (anonymous):

would the equation be y=0x+3?

OpenStudy (anonymous):

yes, slope 0 is correct, but b is wrong, how did you get b = 3?

OpenStudy (anonymous):

guessed

OpenStudy (anonymous):

y = 0x + b, so we substitute the (2, 4) into this, x= 2, y = 4 4 = 0*2 + b 4 = 0 + b 4 = b

OpenStudy (anonymous):

so, then the equation is y=4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how would you figure out b for this problem: (0,5) and (1,3)?

OpenStudy (anonymous):

so, far i have y=1x+b

OpenStudy (anonymous):

wops (5-3)/(0-1) = 2 / (-1) = -2 slope would be -2

OpenStudy (anonymous):

y=-2+5?

OpenStudy (anonymous):

y = -2x + 5

OpenStudy (anonymous):

how would you figure out b for this problem: (0,-1) and (2,3)? This is what I did: (3-(-1)/(2-0)=4/2: =2x -1=2 (0)+b -1=0+b -1=b Equation= 2x+-1

OpenStudy (anonymous):

yup y = 2x - 1

OpenStudy (anonymous):

How would you check your work to see if your right?

OpenStudy (anonymous):

you substitute the other point in y = 2x - 1 substitute (2, 3)

OpenStudy (anonymous):

how would you figure out b for this problem: (0,0) and (2,3)? This is what I did: (3-0)/(2-0)=3/2= 1 1/2 what do you do next?

OpenStudy (anonymous):

3/2 x + b = y stubstitute 0,0 0 + b = 0 b = 0